# Rhaly operators acting on Hardy and Bergman spaces

### v1

@@ 3. Rhaly operators acting on Hardy spaces
$\mathcal{R}_{(\eta)}(f_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}\left(\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n}\right)z^{k}\,=\,\sum_{k=1}^{\infty}\eta_{k}k\left(\frac{1}{kN^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n}\right)z^{k}.$
For $k,N\in\mathbb{N}$, set
(3.4) $\alpha_{k,N}\,=\,\frac{1}{kN^{2-\frac{1}{p}}}\sum_{n=1}^{k}na_{N}^{n},\quad\beta_{k,N}\,=\,\frac{1}{\alpha_{k,N}}.$
Then
(3.5) $\mathcal{R}_{(\eta)}(f_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}k\alpha_{k,N}z^{k},\quad z\in\mathbb{D}.$
Let us point out some properties of the $\alpha_{k,N}$’s and the $\beta_{k,N}$’s.
(3.6) $0\,\leq\,\alpha_{k,N}\,\leq\,\frac{1}{N^{2-\frac{1}{p}}}\frac{1}{k}\sum_{n=1}^{k}n\,=\,\frac{1}{N^{2-\frac{1}{p}}}\frac{k+1}{2}.$
(3.7) $\alpha_{k,N}\geq\frac{1}{N^{2-\frac{1}{p}}}a_{N}^{k}\frac{1}{k}\sum_{n=1}^{k}n\,=\,\frac{a_{N}^{k}}{N^{2-\frac{1}{p}}}\frac{k+1}{2}.$
Using (3.6) and (3.7), we see that there exist two positive constants $C_{1}$ and $C_{2}$ such that
(3.8) $\frac{C_{1}}{N^{1-\frac{1}{p}}}\,\leq\,\alpha_{k,N}\,\leq\,\frac{C_{2}}{N^{1-\frac{1}{p}}},\quad N=,2,3,\dots,\quad N\leq k\leq 2N.$
We have also, for all $k,N$,
$|\alpha_{k+1,N}-\alpha_{k,N}|\,=\,$ $\frac{1}{N^{2-\frac{1}{p}}}\left|\frac{1}{k+1}\sum_{n=1}^{k+1}na_{N}^{n}\,-\,\frac{1}{k}\sum_{n=1}^{k}a_{N}^{n}\right|$
[…]
$\leq\,$ $\frac{3}{2N^{2-\frac{1}{p}}}.$
Using this, (3.6), (3.7), and (3.8), we deduce that
(3.9) $|\beta_{k+1,N}-\beta_{k,N}|\,=\,\frac{|\alpha_{k+1,N}-\alpha_{k,N}|}{|\alpha_{k+1,N}\alpha_{k,N}|}\,\lesssim N^{-\frac{1}{p}},\quad N\leq k\leq 2N,$
and then, we obtain,
(3.10) $\left|\frac{\beta_{k+1,N}-\beta_{k,N}}{\frac{k+1}{N}-\frac{k}{N}}\right|\,=\,N\left|\beta_{k+1,N}-\beta_{k,N}\right|\,\lesssim N^{1-\frac{1}{p}},\quad N\leq k\leq 2N.$
For each $N$, let $\Psi_{N}:\mathbb{R}\rightarrow\mathbb{R}$ be defined as follows:
  • $\Psi_{N}(x)=0$, if $x\leq 0$.
  • $\Psi_{N}(x)=0$, if $x\geq 4$.
  • Between 0 and $4$, the graph of $\Psi_{N}$ is the polygonal whose vertices are the points
    $(0,0),\,\left(\frac{N}{N},\beta_{N,N}\right),\,\left(\frac{N+1}{N},\beta_{N+1,N}\right),\,\left(\frac{N+2}{N},\beta_{N+2,N}\right),\,\dots,\left(\frac{2N}{N},\beta_{2N,N}\right),\,(4,0).$
Using (3.10), we see that $\Psi_{N}$ is a Lipschitz function and that
(3.11) $\mathcal{L}(\Psi_{N})\,\lesssim N^{1-\frac{1}{p}}.$
Taking $\Psi=\Psi_{N}$ and $n=N$ in section 2.3, and using (2.17), we obtain
(3.12) $\left|\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)e^{ik\theta}\right|\,\lesssim N^{1-\frac{1}{p}}\frac{1}{|1-e^{i\theta}|^{2}},\quad 0<\theta<2\pi.$
For each $N$ define $H_{N}$ by
$H_{N}(z)\,=\,\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)z^{k},\quad z\in\mathbb{D}.$
Using (3.12), we deduce that
(3.13) $\|H_{N}\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}.$
Using (2.14), the fact that the $\|\cdot\|_{H^{p}}$ norm increases with $p$, and (3.13), it follows that
(3.14) $\|H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\leq$ $\|H_{N}\|_{H^{1}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\lesssim\|H_{N}\|_{H^{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Set
$G_{(\eta)}(z)=zF^{\prime}_{(\eta)}(z)=\sum_{k=1}^{\infty}k\eta_{k}z^{k},\quad z\in\mathbb{D}.$
Bearing in mind the definitions of $\Psi_{N}$ and $H_{N}$, we see that
$\Delta_{N}(G_{(\eta)})(z)\,=\,\sum_{k=N}^{2N-1}k\eta_{k}z^{k}$
$=\,$ $H_{N}\star\mathcal{R}_{(\eta)}(f_{N})(z)\,-S_{N-1}(H_{N}\star\mathcal{R}_{(\eta)})(f_{N}))(z)\,-\,S_{2N,4N+1}(H_{N}\star\mathcal{R}_{(\eta)})(f_{N}))(z)$
$=\,$ $H_{N}\star\mathcal{R}_{(\eta)}(f_{N})(z)\,-S_{N-1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right)(z)$
$-S_{4N+1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right)(z)+S_{2N-1}\left(H_{N}\star\mathcal{R}_{(\eta)}(f_{N})\right))(z).$
Then, using (3.14) and the Riesz projection theorem, we deduce that
$\|\Delta_{N}(G_{(\eta)})\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Clearly, this implies that
(3.15) $\|\Delta_{N}(F^{\prime}_{(\eta)})\|_{H^{p}}\,\lesssim N^{1-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{p}}.$
Using (3.3), (3.2), (3.15), and Theorem A, we obtain that $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
@@
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$. Since $f_{N}\to 0$, as $N\to\infty$, uniformly in compact subsets of $\mathbb{D}$, using Lemma 3. 7 of [40], we see that $\mathcal{R}_{(\eta)}(f_{N})\to 0$, as $N\to\infty$, in $H^{p}$. This and (3.15) give that
@@
$\|\Delta_{N}(F^{\prime}_{(\eta)})\|_{H^{p}}\,=\,\operatorname{o}\left(N^{1-\frac{1}{p}}\right),\quad\text{as $N\to\infty$}.$
This gives that $F_{(\eta)}\in\lambda^{p}_{1/p}$$\Box$
@@
A result of Flett [13, Theorem 1] (see also Corollary 3. 1 of [29]) asserts that
@@
(iii) If  $1<q<p<\infty$ and $F_{(\eta)}\in\lambda^{q}_{1/q}$, them $\mathcal{R}_{(\eta)}$ is a compact operator from $H^{p}$ into $\mathcal{X}_{q,p}$.
@@
(3.20) $S[f](z)\,=\,zf(z),\quad z\in\mathbb{D},\,\,\,f\in\rm{Hol}(\mathbb{D}).$
Also, let us notice that, for $h\in\rm{Hol}(\mathbb{D})$, we have
@@
We have,
@@
Since $F_{(\eta)}\in\Lambda^{q}_{1/q}$,   $\sup_{0<r<1}\beta_{q}\left(r,F_{(\eta)}\right)<\infty$. Then, using Theorem 5. 11 of [11], it follows that
@@
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers. If $1<q\leq p<\infty$ and $f\in\lambda^{q}_{1/q}$, then $\mathcal{R}_{(\eta)}$ is compact form $H^{p}$ into $\mathcal{X}_{q,p}$.
Proof. Suppose that $1<q\leq p<\infty$ and $f\in\lambda^{p}_{1/p}$. We already know that $\mathcal{R}_{(\eta)}$ is bounded from $H^{p}$ into $\mathcal{X}_{q,p}$. For $N=2,3,4,\dots$ let $\left(\eta^{(N)}\right)\,=\,\{\eta^{(N)}_{n}\}_{n=0}^{\infty}$ be the sequence defined by $\eta^{(N)}_{n}=0$, if $0\leq n\leq N$, and $\eta^{(N)}_{n}=\eta_{n}$, if $n\geq N+1$. Set also $\mathcal{R}_{N}:\,\rm{Hol}(\mathbb{D})\rightarrow\rm{Hol}(\mathbb{D})$ be the operator defined by
@@
4. Rhaly operators acting on Bergman spaces
Theorem 3 follows from the following one.
Theorem 6.
Let $(\eta)\,=\,\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(iii) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$ then $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(iv) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$ then $F_{(\eta)}\in\lambda^{p}_{1/p}$.
Proof. Let us start proving (i) and (ii). So, suppose that $1\leq p<\infty$, $\alpha>-1$, and $F_{(\eta)}\in\Lambda^{p}_{1/p}$. Take $f\in A^{p}_{\alpha}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$). Just as in the proof of Theorem 5, set $g(z)=\frac{f(z)}{1-z}$ ($z\in\mathbb{D}$). Notice that using Hölder’s inequality we have
$M_{1}(r,g)\,\lesssim\frac{1}{(1-r)^{1/p}}M_{p}(r,f).$
Using these facts, (3.23), Theorem 4. 28 of [43], the analogue of (3.21) for the space $A^{p}_{p+\alpha}$, and (2.15), we obtain,
(4.1) $\left\|\mathcal{R}_{(\eta)}(f)\right\|_{A^{p}_{\alpha}}^{p}$ $\,\asymp\,\left|\mathcal{R}_{(\eta)}(f)(0)\right|^{p}\,+\,\int_{\mathbb{D}}(1-|z|)^{p+\alpha}\left|\left[\mathcal{R}_{(\eta)}(f)\right]^{\prime}(z)\right|^{p}dA(z)$
$\lesssim|\eta_{0}|^{p}|a_{0}|^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r^{2},\left[R_{(\eta)}(f)\right]^{\prime}\right)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r^{2},F_{(\eta)}^{\prime}\star g\right)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}M_{p}^{p}\left(r,F_{(\eta)}^{\prime}\right)M_{1}^{p}(r,g)\,dr$
$\lesssim|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{p+\alpha}(1-r)^{1-p}\beta_{p,\frac{1}{p}}^{p}\left(r,F_{(\eta)}\right)\frac{M_{p}^{p}(r,f)}{1-r}\,dr$
$=|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{\alpha}\beta_{p,\frac{1}{p}}^{p}(r,F_{(\eta)})M_{p}^{p}(r,f)\,dr.$
Then, using the fact that $F_{(\eta)}\in\Lambda^{p}_{1/p}$, we obtain that
$\left\|\mathcal{R}_{(\eta)}(f)\right\|_{A^{p}_{\alpha}}^{p}\,\lesssim\,|\eta_{0}|^{p}\|f\|_{A^{p}_{\alpha}}^{p}\,+\,\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f)\,dr\,\lesssim\|f\|_{A^{p}_{\alpha}}^{p}.$
Hence, we have proved (i).
Suppose now that $1<p<\infty$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$. For $N=2,3,4,\dots$ let $\mathcal{R}_{N}$ be defined as in (3.27). Using (4.1) and arguing as in the proof of Proposition 2, we deduce, for $f\in A^{p}_{\alpha}$,
$\left\|\left(\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N}\right)(f)\right\|_{A^{p}_{\alpha}}^{p}\lesssim$ $\int_{0}^{1}\beta^{p}_{p,\frac{1}{p}}\left(r,F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)(1-r)^{\alpha}M_{p}^{p}(r,f)^{p}\,dr$
$\lesssim$ $\,\beta^{p}_{p,\frac{1}{p}}\left(F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)\|f\|_{A^{p}_{\alpha}}^{p}.$
Using Proposition 1 we deduce that $\{\mathcal{R}_{(\eta)}-\mathcal{R}_{N}\}\,\rightarrow\,0$, as $N\to\infty$, in the norm of $\mathcal{B}(A^{p}_{\alpha})$. Since the operators $\mathcal{R}_{N}:A^{p}_{\alpha}\rightarrow A^{p}_{\alpha}$ are finite rank operators, it follows that $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$. Thus, (ii) is proved.
The proof of parts (iii) and (iv) is closely related to that of the necessity statements in Theorems 1 and 2.
Suppose that $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$.
For $N=2,3,4,\dots$, set $a_{N}=1-\frac{1}{N}$ and
$g_{N}(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,N^{\frac{\alpha+1}{p}}f_{N}(z),\quad z\in\mathbb{D},$
where $f_{N}$ is defined in (3.1). Since $\alpha<2p-2$, Lemma 3. 10 of [43] shows that $g_{N}\in A^{p}_{\alpha}$ for all $N$ and $\|g_{N}\|_{A^{p}_{\alpha}}\asymp 1$. For $k,N\in\mathbb{N}$, set
(4.2) $\gamma_{k,N}\,=\,N^{\frac{1+\alpha}{p}}\alpha_{k,N},\quad\delta_{k,N}\,=\,\frac{1}{\gamma_{k,N}},$
where $\alpha_{k,N}$ and $\beta_{k,N}$ are defined in (3.4). Using (3.5), we see that
(4.3) $\mathcal{R}_{(\eta)}(g_{N})(z)\,=\,\sum_{k=1}^{\infty}\eta_{k}k\gamma_{k,N}z^{k},\quad z\in\mathbb{D}.$
We have

### v2

@@ 4. Necessary conditions
$\mathcal{R}_{(\eta)}(f_{N,p})(z)\,=\,\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{\infty}\eta_{n}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)z^{n},\quad z\in\mathbb{D}.$
Then it follows that
(4.6) $\Delta_{N}(F_{(\eta)})(z)\,=\,\sum_{n=N}^{2N-1}\eta_{n}z^{n}\,=\,N^{2-\frac{1}{p}}\mathcal{R}_{(\eta)}(f_{N,p})\star\varphi_{N}(z),\quad z\in\mathbb{D},\,\,N\geq 2.$
Using Proposition 2, we obtain
(4.7) $\|\Delta_{N}(F_{(\eta)})\|_{H^{p}}\,$ $\lesssim\,N^{2-\frac{1}{p}}\|R_{(\eta)}(f_{N,p})\|_{H^{p}}\|\varphi_{N}\|_{H^{1}}\,\lesssim\,N^{-\frac{1}{p}}\log N\|R_{(\eta)}(f_{N,p})\|_{H^{p}}.$
Then, using Theorem 2 of [21], (2.6), and (4.5), this yields that $F_{(\eta)}$ satisfies (1.2).
@@
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$. Since $f_{N,p}\to 0$, as $N\to\infty$, uniformly in compact subsets of $\mathbb{D}$, using Lemma 3. 7 of [42], we see that $\mathcal{R}_{(\eta)}(f_{N,p})\to 0$, as $N\to\infty$, in $H^{p}$. This and (4.7) give that
@@
$\|\Delta_{N}(F_{(\eta)})\|_{H^{p}}\,=\,\operatorname{o}\left(\left(N^{-\frac{1}{p}}\right)\log N\right),\quad\text{as $N\to\infty$}.$
This and the “little oh”  version of Theorem 2 of [21] give that $F_{(\eta)}$ satisfies (1.3) $\Box$
Proof ofTheorem 3 (iii) and Theorem 3 (iv). Assume that $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$.
For $N=2,3,4,\dots$, set
$g_{N,p}(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,N^{\frac{\alpha+1}{p}}f_{N,p}(z),\quad z\in\mathbb{D},$
where $f_{N,p}$ is defined in (4.3). We have that $g_{N,p}\in A^{p}_{\alpha}$ for all $N$ and, since $\alpha<2p-2$, Lemma 3. 10 of [45] shows that and $\|g_{N,p}\|_{A^{p}_{\alpha}}\asymp 1$. Using (4.6), we see that
$\Delta_{N}(F_{(\eta)})\,=\,N^{\frac{-(\alpha+1)}{p}}N^{2-\frac{1}{p}}\mathcal{R}_{(\eta)}(g_{N.p})\star\varphi_{N}.$
This, Lemma 2, and Proposition 2 give
(4.8) $\|\Delta_{N}(F_{(\eta)})\|_{A^{p}_{\alpha}}\,\lesssim N^{\frac{-(\alpha+1)}{p}}N^{2-\frac{1}{p}}\|\mathcal{R}_{(\eta)}(g_{N.p})\|_{A^{p}_{\alpha}}\frac{\log N}{N^{2}}\,=\,N^{\frac{-(\alpha+1)}{p}}N^{-\frac{1}{p}}\log N\|\mathcal{R}_{(\eta)}(g_{N.p})\|_{A^{p}_{\alpha}}.$
Using Lemma 3. 1 of [31] it follows easily that $\|\Delta_{N}\left(F_{(\eta)}\right)\|_{H^{p}}\asymp\,N^{\frac{1+\alpha}{p}}\|\Delta_{N}\left(F_{(\eta)}\right)\|_{A^{p}_{\alpha}}$. Then it follows that $\|\Delta_{N}\left(F_{(\eta)}\right)\|_{H^{p}}\lesssim N^{-\frac{1}{p}}\log N$, and, hence, $F_{(\eta)}$ satisfies (1.2).
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$. Since $g_{N,p}\,\rightarrow\,0$, as $N\to\infty$, uniformly in compact subsets of $\mathbb{D}$, it follows that $\mathcal{R}_{(\eta)}(g_{N,p})\,\rightarrow\,0$, as $N\to\infty$, in $A^{p}_{\alpha}$. This leads to $\|\Delta_{N}\left(F_{(\eta)}\right)\|_{A^{p}_{\alpha}}\,=\operatorname{o}\left(N^{\frac{-(\alpha+1)}{p}}N^{-\frac{1}{p}}\log N\right)$ and to $\|\Delta_{N}\left(F_{(\eta)}\right)\|_{H^{p}}=\,\operatorname{o}\left(N^{-\frac{1}{p}}\log N\right)$. Hence, $F_{(\eta)}$ satisfies (1.3). $\Box$
Proof ofTheorem 4 (iii) and Theorem 4 (iv). Assume that $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(\mathcal{D}^{p}_{\alpha})$.
For $N=2,3,4,\dots$, set
$h_{N,p}(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\frac{a_{N}z}{1-a_{N}z}\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\sum_{n=1}^{\infty}a_{N}^{n}z^{n},\quad z\in\mathbb{D}.$
Then $h_{N,p}\in\mathcal{D}^{p}_{\alpha}$ for all $N$ and, using the fact that $2p-\alpha-2>0$ and Lemma 3. 10 of [45], we see that
$\|h_{N,p}\|_{\mathcal{D}^{p}_{\alpha}}^{p}\,\asymp\,\frac{1}{N^{2p-\alpha-2}}\int_{\mathbb{D}}\frac{(1-|z|)^{\alpha}}{|1-a_{N}z|^{2p}}\,dA(z)\,\asymp\,1.$
This implies that $\|\mathcal{R}_{(\eta)}(h_{N,p})\|_{\mathcal{D}^{p}_{\alpha}}\lesssim 1$.
(4.9) $S[f](z)\,=\,zf(z),\quad z\in\mathbb{D},\,\,\,f\in\rm{Hol}(\mathbb{D}).$
We have
$\mathcal{R}_{(\eta)}(h_{N,p})(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\sum_{n=1}^{\infty}\eta_{n}\left(\sum_{k=1}^{n}a_{N}^{k}\right)z^{n},\quad z\in\mathbb{D},$
and, hence,
$S\left(R_{(\eta)}(h_{N,p})^{\prime}\right)(z)\,=\,\frac{1}{N^{2-\frac{\alpha+2}{p}}}\sum_{n=1}^{\infty}n\eta_{n}\left(\sum_{k=1}^{n}a_{N}^{k}\right)z^{n},\quad z\in\mathbb{D}.$
Then it follows that
(4.10) $\Delta_{N}\left(S(F_{\eta}^{\prime})\right)\,=\,N^{2-\frac{\alpha+2}{p}}S\left(R_{(\eta)}(h_{N,p})^{\prime}\right)\star\chi_{N},$
where,
$\chi_{N}(z)\,=\,\sum_{n=N}^{2N-1}\frac{1}{\sum_{k=1}^{n}a_{N}^{k}}z^{n},\quad z\in\mathbb{D}.$
Arguing as in the proof of Proposition 4.1, we deduce that
(4.11) $\|\chi_{n}\|_{H^{1}}\,\asymp\,\frac{\log N}{N}.$
Using this, (4.10), and Lemma 2, we obtain
$\|\Delta_{N}(F_{(\eta)}^{\prime})\|_{A^{p}_{\alpha}}\,\lesssim N^{1-\frac{\alpha+2}{p}}\log N\|\mathcal{R}_{(\eta)}(h_{N,p})\|_{\mathcal{D}^{p}_{\alpha}}.$
Then Lemma 3. 1 of [31] implies that
$\|\Delta_{N}(F_{(\eta)}^{\prime})\|_{H^{p}}\,\lesssim N^{1-\frac{1}{p}}\log N\|\mathcal{R}_{(\eta)}(h_{N,p})\|_{\mathcal{D}^{p}_{\alpha}}\,\lesssim N^{1-\frac{1}{p}}\log N$
and this gives that $F_{(\eta)}$ satisfies (1.2).
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(\mathcal{D}^{p}_{\alpha})$. Since $h_{N,p}\,\rightarrow\,0$, as $N\to\infty$, uniformly in compact subsets of $\mathbb{D}$, it follows that $\mathcal{R}_{(\eta)}(h_{N,p})\,\rightarrow\,0$, as $N\to\infty$, in $\mathcal{D}^{p}_{\alpha}$. This leads to
$\|\Delta_{N}(F_{(\eta)}^{\prime})\|_{A^{p}_{\alpha}}\,=\,\operatorname{o}\left(N^{1-\frac{\alpha+2}{p}}\log N\right)$
and to
$\|\Delta_{N}(F_{(\eta)}^{\prime})\|_{H^{p}}\,=\,\operatorname{o}\left(N^{1-\frac{1}{p}}\log N\right),$
which implies that $F_{(\eta)}$ satisfies (1.3). $\Box$
5. Sufficient conditions
@@
A result of Flett [14, Theorem 1] (see also Corollary 3. 1 of [31]) asserts that
@@
(iii) If  $1<q<p<\infty$ and $F_{(\eta)}\in\lambda^{q}_{1/q}$, then $\mathcal{R}_{(\eta)}$ is a compact operator from $H^{p}$ into $\mathcal{X}_{q,p}$.
@@
Let us notice that, for $h\in\rm{Hol}(\mathbb{D})$, we have
@@
With $S$ denoting the shift operator as above, we have,
@@
Since $F_{(\eta)}\in\Lambda^{q}_{1/q}$,   $\sup_{0<r<1}\beta_{q,\frac{1}{q}}\left(r,F_{(\eta)}\right)<\infty$. Then, using Theorem 5. 11 of [12], it follows that
@@
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers. If $1<q\leq p<\infty$ and $F_{(\eta)}\in\lambda^{q}_{1/q}$, then $\mathcal{R}_{(\eta)}$ is compact form $H^{p}$ into $\mathcal{X}_{q,p}$.
Proof. Suppose that $1<q\leq p<\infty$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$. We already know that $\mathcal{R}_{(\eta)}$ is bounded from $H^{p}$ into $\mathcal{X}_{q,p}$. For $N=2,3,4,\dots$ let $\left(\eta^{(N)}\right)\,=\,\{\eta^{(N)}_{n}\}_{n=0}^{\infty}$ be the sequence defined by $\eta^{(N)}_{n}=0$, if $0\leq n\leq N$, and $\eta^{(N)}_{n}=\eta_{n}$, if $n\geq N+1$. Set also $\mathcal{R}_{N}:\,\rm{Hol}(\mathbb{D})\rightarrow\rm{Hol}(\mathbb{D})$ be the operator defined by
@@
Since, for $\alpha>-1$ and $p>0$, $A^{p}_{\alpha}=\mathcal{D}^{p}_{\alpha+p}$ (see Theorem 4. 28 of [45]), parts (i) and (ii) of Theorem 3 follow from parts (i) and (ii) of Theorem 4.
Proof ofTheorem 4 (i) and Theorem 4 (ii). Assume that $1\leq p<\infty$, $\alpha>p-2$, and $F_{(\eta)}\in\Lambda^{p}_{1/p}$. Using Theorem 4. 28 of [45], we have to prove that, for $f\in\mathcal{D}^{p}_{\alpha}$,
(5.13) $\int_{\mathbb{D}}(1-|z|)^{\alpha+p}|\mathcal{R}_{(\eta)}(f)^{\prime\prime}(z)|^{p}\,dA(z)\,\lesssim\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
This is equivalent to
(5.14) $\int_{0}^{1}(1-r)^{\alpha+p}M_{p}^{p}(r^{4},\mathcal{R}(f)^{\prime\prime})\,dr\,\lesssim\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
So, take $f\in\mathcal{D}^{p}_{\alpha}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$). Just as in the proof of Theorem 7, set $g(z)=\frac{f(z)}{1-z}$ ($z\in\mathbb{D}$). Bearing in mind (5.7) and the continuity of the shift operator on $H^{p}$, we see that (5.14) will follow from
(5.15) $\int_{0}^{1}(1-r)^{\alpha+p}M_{p}^{p}(r^{4},F_{(\eta)}^{\prime}\star g^{\prime})\,dr\,\lesssim\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
Notice that
$g^{\prime}(z)\,=\,\varphi(z)\,+\,\phi(z),\quad z\in\mathbb{D},$
with
$\varphi(z)\,=\,\frac{f^{\prime}(z)}{1-z},\quad\phi(z)\,=\,\frac{f(z)}{(1-z)^{2}},\quad z\in\mathbb{D}.$
Using this, (2.17), and the fact that $F_{(\eta)}\in\Lambda^{p}_{1/p}$, we see that
(5.16) $\int_{0}^{1}(1-r)^{\alpha+p}M_{p}^{p}(r^{4},F_{(\eta)}^{\prime}\star g^{\prime})\,dr\lesssim\int_{0}^{1}(1-r)^{\alpha+p}M_{p}^{p}(r^{2},F_{(\eta)}^{\prime})M_{1}^{p}(r^{2},g^{\prime})\,dr$
$\lesssim\beta_{p,\frac{1}{p}}^{p}(F_{(\eta)})\left[\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\varphi)\,dr+\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\phi)\,dr\right]$
$\lesssim\beta_{p,\frac{1}{p}}^{p}(F_{(\eta)})\left[\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r,\varphi)\,dr+\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\phi)\,dr\right].$
Using Hölder’s inequality, we see that
$M_{1}(r,\varphi)=\frac{1}{2\pi}\int_{0}^{2\pi}\frac{|f^{\prime}(re^{i\theta})|}{|1-re^{i\theta}|}\,d\theta\lesssim\frac{1}{(1-r)^{1/p}}M_{p}(r,f^{\prime}).$
This implies that
(5.17) $\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r,\varphi)\,dr$ $\,\lesssim\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f^{\prime})\,dr\,\lesssim\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
Take now $\gamma>1$ such that $\alpha>p-2+\frac{1}{\gamma}$. Let $\sigma$ be the exponent conjugate to $\gamma p$, that is, $\frac{1}{\gamma p}+\frac{1}{\sigma}=1$. Using Hölder’s inequality again, we see that